OpenStudy (anonymous):

what is the area of a garden whose perimeter is 36 and length is 4? help me

OpenStudy (apoorvk):

Is this a square-shaped garden?

hartnn (hartnn):

garden shape ?

OpenStudy (anonymous):

rectangular ..sorry

OpenStudy (anonymous):

Perimeter and Area are both based on length and width, try this P = 2l + 2w A =lw

OpenStudy (anonymous):

perimeter = 2 length + 2 breadth length = 4 solve them as u know the perimeter

OpenStudy (anonymous):

okay 36= 2(length+width ) 36= 2(4+ w) now?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

find width

hartnn (hartnn):

so divide by 2 18=4+w w= 14 ok?

OpenStudy (anonymous):

solve that eqn to find the width..

OpenStudy (anonymous):

P = 2l +2w and you know l=4 and P=36, then solve for width and use it to find area!

OpenStudy (anonymous):

thanks....and now ?

hartnn (hartnn):

now area = l*w = 4*14 = ??

OpenStudy (anonymous):

okay is it 56?

hartnn (hartnn):

yup :)

OpenStudy (anonymous):

yaa...

OpenStudy (anonymous):

Thank you guys (: